Skip to content
Back to Top

Site Report: Thane Bunker Fuel Tanks


Site Name: Thane Bunker Fuel Tanks
Address: Head of Thane Road, Juneau, AK 99801
File Number: 1513.38.018
Hazard ID: 899
Status: Cleanup Complete
Staff: ,
Latitude: 58.294900
Longitude: -134.397200
Horizontal Datum:NAD83

We make every effort to ensure the data presented here is accurate based on the best available information currently on file with DEC. It is therefore subject to change as new information becomes available. We recommend contacting the assigned project staff prior to making decisions based on this information.

Problems/Comments

Remnants of bulk storage tanks containing old bunker fuel. Tanks have leaked in the past and pose a potential threat to nearby Gastineau Channel. Numerous efforts to get responsible party to contain and remediate site have failed. Requested funds for remediation 8/1990, approved. Plan finalized by ADEC to remove oil 1/1990, funds re-requested from new administration, per 1/28/1991 memorandum. 1996: New project manager -- Bob Mattson of the Prevention and Emergency Response Program. Plans a preliminary assessment. Contaminated Sites will maintain an advisory role.

Action Information

Action Date Action Description DEC Staff
8/26/1993 Site Added to Database No Longer Assigned
11/10/1994 Site Ranked Using the AHRM Initial ranking. No Longer Assigned
12/21/1994 Update or Other Action Corrective action by JDO and RP agreed to and pending. No Longer Assigned
2/26/1996 Update or Other Action This site is currently being handled under PERP by Bob Mattson of the SE Area Response Team. A preliminary assessment out of this office is forthcoming, and the site may become a state-lead cleanup. Sally Schlichting
11/5/1997 Update or Other Action Site transferred from the PERP program back to CS. CS plans no immediate action for the present. VCP letter sent to Chuck Keen. Sally Schlichting
10/15/1999 Update or Other Action EPA contacted Mr. Keen during the last two years regarding a current SPCC Plan. Do not know outcome of the communications. Bill Janes
2/20/2002 Update or Other Action Project tickler update - No action at this time. Bill Janes
1/7/2004 Meeting or Teleconference Held Landowner reps. came to office to discuss possible remediation options. Informed they need to bring on a qualified third party consultant. Bill Janes
4/21/2005 Update or Other Action Comments send to DMC Tech on the draft remedial workplan received early April. Bunker C and sludge to be incinerated on-site. Contaminated soil to be bio-treated on-site. Bill Janes
5/30/2005 Update or Other Action Contractor installing site control perimeter fence this week. Bill Janes
6/3/2005 Update or Other Action Telecon with Peter Freer at CBJ briefing him on project. Bill Janes
6/6/2005 Update or Other Action Additonal comments sent to DMC Tech on the revised draft cleanup workplan. Bill Janes
8/12/2005 Cleanup Level(s) Approved Method 2 migration to groundwater Bill Janes
8/12/2005 Cleanup Plan Approved All comments resolved Bill Janes
8/31/2005 Update or Other Action Pre-cleanup site visit and photos Bill Janes
9/16/2005 Update or Other Action Pre-cleanup site visit to discuss water issues. Excess water in the tanks will be treated through wastewater treatment plant before release into Gastineau Channel. Bruce Wanstall
9/19/2005 Site Visit Site visit to talk to press Bill Janes
11/4/2005 Site Visit Site visit with Larry Dietrick and to meet with KTOO reporter. Burners have arrived and have started processing the bunker-C on a pilot basis. Oil and debris are placed in a large holding tank where it is heated to 180 degrees. It is then run through a grinder and the debris is reduced to 1/4 inch. Bill Janes
11/29/2005 Update or Other Action Open burn permit denied by Air Quality Division. Bill Janes
12/5/2005 Update or Other Action Project put on hold until next spring. Interim report filed by contractor. Project estimated 75% complete. Technical changes to the burners underway. The atomizer fan will be twice as large. Pump capacity will be increased. Pumps will be able to pump at 200 degrees versus 140. Hot air will be re-circulated to increase combustion efficiency. Bill Janes
2/21/2007 Potentially Responsible Party/State Interest Letter Letters sent to AJ Land Company and AJ Industries. Meeting proposed April 9. Bill Janes
3/29/2007 Update or Other Action Meeting with Asst. AG Ruth Heese regarding PRP letter response. It appears that AJ Industries went bankrupt in the 1990s. The attorney for AJ Land Company responded to the letter and stated AJ Land Company never had an interest in the property. Ruth is sending a follow-up letter with documentation and suggesting a late April or early May meeting date. The letter will also explain that the state is moving ahead to obtain an accurate cost estimate for removal of the sludge to an out of state facility. Bill Janes
4/27/2007 GIS Position Updated NAD 83 using CS web map air photo Bill Janes
4/27/2007 Site Visit Site visit with Shannon and Wilson reps. to obtain accurate cost estimate for removing oil and debris from all tanks. Bill Janes
5/4/2007 Update or Other Action Second PRP letter sent to AJ Land Company. AJ Industries apparently bankrupt. May 29 conference call proposed. Bill Janes
6/12/2007 Update or Other Action PRP meeting. PRP's will discuss options and come back to DEC in two months. If they do not have a viable plan we will consider initiating a state lead cleanup. Bill Janes
8/10/2007 Update or Other Action Follow-up PRP meeting. One PRP has a draft plan on the table and is expecting another next week. Telecon status update scheduled for Friday, August 17. Bill Janes
8/17/2007 Update or Other Action Follow-up PRP meeting to review the second proposal. DEC asked to provide initial comments on both proposals. Bill Janes
9/28/2007 Update or Other Action Draft removal action work plan received this date. Bill Janes
10/1/2007 Interim Removal Action Approved Comments on draft work plan sent to Bureau Veritas. Bill Janes
10/4/2007 Update or Other Action Bureau Veritas arrived on site and is doing preparation work for the cleanup. Bill Janes
10/17/2007 Site Visit Site visit conducted. 65,000 gallons water pumped into the city sewer system to date. Diatomaceous earth is now being added to the sludge and the first four or five containers should be filled by Friday for shipment to Seattle. A few barrels of product that has been "cut" with diesel and used at the Bicknell asphalt plant is awaiting testing at the plant to determine if it is usable. Bill Janes
10/19/2007 Site Visit Site visit conducted. Bill Janes
10/26/2007 Site Visit Site visit conducted. Bill Janes
10/26/2007 Update or Other Action Approval letter sent to Bicknell to accept approximately 26,000 gallons of oil cut with 10-20% diesel for use in their asphalt plant as kiln fuel. Bill Janes
11/16/2007 Site Visit Site visit conducted. Bill Janes
11/19/2007 Site Visit Site visit conducted. Bill Janes
11/21/2007 Update or Other Action Situation Report received from Bob Fultz, Ketchikan DEC, ragarding bunker oil that leaked from one of the containers. Approximately 200 gallons of oily waste escaped and required a response action by the Coast Guard and SEAPRO. Contaminated material removed frmo the damage container and re-loaded into haz-mat bags. The bags will be loaded into another container for continued shipment south. Bill Janes
11/28/2007 Update or Other Action Site visit conducted. Bill Janes
11/28/2007 Update or Other Action Site visit conducted. Bill Janes
12/4/2007 Update or Other Action Site inspection. North tank has been dismantled and the scrap metal is being loaded for shipment south. A small soil stockpile has been constructed in the footprint of the former tank. About 10 cubic yards of obviously stained soil is being placed on a liner and will be fully wrapped to avoid contact with water. Bill Janes
12/6/2007 Update or Other Action Site visit conducted. Bill Janes
12/11/2007 Update or Other Action Final site inspection conducted. Bill Janes
2/26/2008 Exposure Tracking Model Ranking Initial ranking with ETM completed. Bill Janes
3/20/2008 Update or Other Action Interim Removal Action Completion Report received. Bill Janes
4/23/2008 Update or Other Action Site visit with Steve Haavig of Carson Dorn to scope the next phase of the cleanup project. Workplan expected within a couple of weeks. Pockets of residual heavy oil remain along the sidewalls. Bill Janes
5/1/2008 Update or Other Action Interim Removal Action project closed out and letter sent to PRPs. Bill Janes
5/23/2008 Update or Other Action Site visit with Tom Hanna. Bill Janes
6/10/2008 Site Visit Site visit. Tom Hanna from SE Mgmt Services is overseeing the excavation of contaminated soil along the tank sidewalls. On-site bioremediation will be performed over the summer. Bill Janes
6/10/2008 Cleanup Plan Approved Email approval sent this date. On-site bioremediation proposed. No liner required; bedrock very shallow and additional contamination in the footprint of the tank will be minimal as a result of the bio-treatment. Bill Janes
6/10/2008 Cleanup Level(s) Approved Method 2 ingestion for DRO, RRO and PAHs. Email approval sent this date. Bill Janes
7/8/2008 Update or Other Action Telecon with consultant on-site. He estimates there will be about 500 cy of soil that will need treatment in the former south tank footprint. Biotreatment in the north tank expected to be finished in about a month. Bill Janes
7/28/2008 Report or Workplan Review - Other Site progress report and cleanup plan revision received. Contaminated soil above the DRO and RRO human health levels is thought to be present across about 75% of the former south tank location at a depth of 8-18". A plastic-lined water containment sump has been constructed The upslope drainage ditch has been excavated to allow flow into the sump. The deepened drainage ditch allowed the first 8-12" of contaminated soil to removed over about 1/2 the area and stockpiled A biocell will be constructed on-site (no bottom liner) and the contaminated soil will be spread in an 18" layer. Urea and fertilizer will be added and tilling will occur every two weeks. The parking area adjacent to the north tank will be evaluated when the south tank cleanup is completed. Bill Janes
8/22/2008 Site Visit South tank soil excavation underway for on-site treatment. About 400 cubic yards will be bioremediated. Bill Janes
11/9/2009 Update or Other Action Site visit with Tom Hanna and Bill Janes. It was determine that the stockpile does not need to be covered for the winter based on oberservations indicating no apparent contamination visible at the site and assumed natural attenuation that has occurred since the summer of 2008. Denise Elston
8/7/2012 Report or Workplan Review - Other The 2008-09 cleanup report was delivered by Southeast Management Services. The report was evaluated by DEC staff, Denise Elston, on the cleanup activity that accurred at the site before the Alaska Trams Limited Partnership, ran out of funding and cleanup activity ceased. Denise Elston
8/10/2012 Site Visit DEC staff, Denise Elston and Bruce Wanstall performed a site visit to evaluate the current condition of the site. Site conditions were recorded. The bioremediation cells that were initially set up in 2009 will need further characterization data to confirm addition remediation activity necessary at the site. Denise Elston
11/1/2012 Update or Other Action CSP staff Denise Elston and Sally Schlichting met with Mr. Corbus and consulant to discuss the cleanup plan for the site. Bioremediation efforts will resume in the spring 2013. Denise Elston
3/12/2013 Report or Workplan Review - Other Conditional approval of the Corrective Action Plan submitted by Nortech Denise Elston
4/28/2013 Update or Other Action I received a request form the consultant that the RP would like to sample the bioremediation cells at this time; before fertilizer addition takes place in conjunction with active remediation methods, i.e. tilling. The department approved pre-treatment sampling based on the CAP sampling methodology received in February 2013 that stated MI sampling would be performed. Denise Elston
5/1/2013 Site Visit Contaminated Site staff performed a site visit to observe multi-increment sampling methods as conducted by the consultant. CS staff did not observe work in progress, however' we continues to walk around the site and investigate the site based on physical characteristics. These observations were documented and I will follow up with the consultant. Denise Elston
9/26/2013 Report or Workplan Review - Other Denise Elston
10/11/2013 Site Visit ADEC staffs performed a site visit for a current assessment of the site conditions after MI sampling completed this summer. Denise Elston
12/4/2013 Cleanup Complete Determination Issued The Alaska Department of Environmental Conservation, Contaminated Sites Program (DEC) has completed a review of the environmental records associated with the Thane Bunker Fuel Tanks. After years of regulatory oversight and environmental cleanup and remediation efforts, the DEC has determined that the contaminant concentrations remaining on site do not pose an unacceptable risk to human health or the environment, and this site will be closed. This decision is based on the Thane Bunker Fuel Tanks Contaminated Site administrative record, which is located in the offices of the DEC in Juneau, Alaska. This letter summarizes the decision process used to determine the environmental status of this site and provides a summary of the regulatory issues considered in the Cleanup Complete Determination. A copy of the decision letter can be found in the DEC database for your review. Denise Elston
2/19/2014 Update or Other Action Contacted by ERM consultants out of Anchorage to discuss the transfer of assets from AJT properties to Avista Corp. Sally Schlichting

Contaminant Information

Name Level Description Media Comments
RRO Between Method 2 Migration to Groundwater and Human Health/Ingestion/Inhalation Soil The Multi-Increment soil sampling for RRO determined that the 95th UCL was calculated to be 2787 mg/kg
DRO Between Method 2 Migration to Groundwater and Human Health/Ingestion/Inhalation Soil The Multi-Increment soil sampling for DRO determined that the 95th UCL was calculated to be 1792 mg/kg

Control Type

Type Details
No ICs Required

Requirements

Description Details
Advance approval required to transport soil or groundwater off-site.

Missing Location Data

Because the GIS location data for this site has not been entered, we cannot direct you to its position on the map. Click "Continue" to proceed to the Contaminated Sites Web Map or "Close" to return to the site report.
Continue     Close